OpenStudy (anonymous):

A dimensionless quantity (a) never has a unit, (b) always has a unit, (c) may have a unit, (d) does not exist.

OpenStudy (radar):

I had one of those once, and it did not have a unit.

OpenStudy (anonymous):

What was it then?

OpenStudy (radar):

What is the unit of Pi?

OpenStudy (anonymous):

No unit

OpenStudy (anonymous):

what a unit of 25ºC

OpenStudy (radar):

The unit is Celcius degree @myko

OpenStudy (radar):

Can be converted to Fahrenheit units by a sharp student.

OpenStudy (anonymous):

so the answer is (c) may have a unit. Other definition for dimensionless quantity is scalar quantity

OpenStudy (anonymous):

Yeah But give an example with a unit and without a unit!

OpenStudy (anonymous):

i just did. And @radar gave with no units

OpenStudy (anonymous):

@radar also can convert it to Fahrenheit if you whant

OpenStudy (anonymous):

I need Examples One is temperature with a UNiT! So without unit?

OpenStudy (radar):

In my humble opinion, the best answer would be "a"

OpenStudy (anonymous):

ohh why!?

OpenStudy (radar):

Pi has no unit, just a value and an irrational value at that.

OpenStudy (anonymous):

I desagree: http://en.wikipedia.org/wiki/Dimensionless_quantity

OpenStudy (anonymous):

So why is it not unitless?

OpenStudy (radar):

Plancks constant is just a value, no unit, it is just used to convert things that have a dimehnsion, ft, in. grams, degrees, lbs, etc. these are dimensionsal units.

OpenStudy (anonymous):

oh yes!

OpenStudy (radar):

Review that link provided by @rkparth5770 and make a decision.

OpenStudy (anonymous):

What?! i did not provide any

OpenStudy (radar):

Thanks @rkparth5770 for providing additional info on this "dimensionless" subject.

OpenStudy (anonymous):

Welcum @radar

OpenStudy (callisto):

Hmmm... in my physics book, Planck's constant is 6.626...x10^(-34) Js , which is with a unit :S

OpenStudy (radar):

Just picked that out of the dark @Callisto, didn't realize that Plalnck's constant was in some kind of unit. Should of stuck with Pi lol.

OpenStudy (unklerhaukus):

in SI units planks constant certainly does have dimensions, , that is why Max Plank devise his own set of units that made it, along with other constants like the gravitational constant and the speed of light dimensionless. i dont like the wording of the options, i would phrase the answer as " a dimensionless quantities has units that cancel out" for example \[\pi=\frac{C[{l}]}{d[l]}=\frac{C[\text{cm}]}{d[\text{cm}]}=\frac{C\cancel{[\text{cm}]}}{d\cancel{[\text{cm}]}}=\frac{C}{d}\]

OpenStudy (radar):

Yes indeed Plancks constant does have units. I looked up (Google): Planck's constant = 6.626068 × 10^-34 m^2 kg / s Thanks UnkleRhaukus for the review.

OpenStudy (radar):

Yes indeed Plancks constant does have units. I looked up (Google): Planck's constant = 6.626068 × 10^-34 m^2 kg / s Thanks UnkleRhaukus for the review.

OpenStudy (radar):

Yes indeed Plancks constant does have units. I looked up (Google): Planck's constant = 6.626068 × 10^-34 m^2 kg / s Thanks UnkleRhaukus for the review.